Shift report 20030512 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon May 12 2003 - 07:57:55 EDT

  • Next message: Apache: "Shift report 20030512 08:00-16:00"
    ------------------------------------------------------------
    Date: Mon May 12 07:57:54 2003, Shift: 00:00-08:00
    Supervisor: sjs, with: 
    Log book#14 pages 107 - 107
    
    Shift Summary:
    ------------------------------------------------------------
    Very little accomplished during shift.  The yellow beam 
    polarization drops to about 10% after the rotators, leading 
    to the beam being dumped almost immediately after the beginning of a store.  The first half of the shift had a controlled access to address a problem with the "X arc magnet". 
    Two short data runs were taken while tuning was still taking place:
    Run 8910 (trig 2: 63 counts)
    Run 8913 (trig 2: 41 counts)
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon May 12 2003 - 08:02:35 EDT